# Debiasing Watermarks for Large Language Models via Maximal Coupling

### v1

@@ 1 Introduction
The field of artificial intelligence (AI) has seen rapid growth in recent years, with large language models (LLMs) playing a central role in many applications. These models, such as OpenAI’s ChatGPT and GPT-4 (Achiam et al., 2023), are capable of understanding and generating human-like text. However, the potential for misuse has raised concerns about the trustworthiness of AI-generated content, particularly in the context of disinformation (Hsu and Thompson, 2023), fraud, social engineering, and cybercrime (Europol, 2023). To mitigate these risks, President Biden issued an executive order on Safe, Secure, and Trustworthy Artificial Intelligence, which emphasized the need for standards to detect AI-generated content and authenticate official content, setting a precedent for transparency and trust in digital communications (Biden, 2023). In particular, the Department of Commerce identified watermarking as a pivotal means to clearly label AI-generated content.
@@
The earliest watermarking scheme, the Gumbel-max watermark, was introduced by Aaronson and Kirchner (2023). This method selects the next token $w_{t}$ using the decoder $w_{t}={\mathcal{S}}(\bm{P}_{t},\xi_{t}):=\operatorname{argmax}_{w\in{\mathcal{W}}}\log\frac{U_{w}}{P_{w}}$, assuming $\xi_{t}=(U_{1},\ldots,U_{|{\mathcal{W}}|})$ is a random vector consisting of $|{\mathcal{W}}|$ i.i.d. copies of $U[0,1]$, serving as part of the watermark key. While this watermarked decoder theoretically preserves the original next-token distribution when $\xi_{t}$ is truly random (Li et al., 2024) and achieves high detection efficiency, this decoder can generate repetitive text and significantly degrade text quality in practice (Christ et al., 2023; Kuditipudi et al., 2023). The main reason is that, for existing implementations, the pseudorandom variable $\xi_{t}$ depends deterministically on the prior $k$ tokens $w_{(t-k)},\ldots,w_{(t-1)}$. As a result, when conditioned on previous tokens, the decoder ${\mathcal{S}}$ deterministically outputs the next token, which can lead to sub-optimal text generation (Holtzman et al., 2020).
Another influential watermarking scheme, the green/red list watermark, employs a stochastic decoder. First introduced by Kirchenbauer et al. (2023a), this approach has recently gained popularity (Zhao et al., 2023; Fernandez et al., 2023; Kirchenbauer et al., 2023b; Hu et al., 2023; Wu et al., 2023). It first partitions the token vocabulary ${\mathcal{W}}$ into the green list ${\mathcal{G}}$ and the red list ${\mathcal{R}}={\mathcal{W}}\setminus{\mathcal{G}}$ (pseudo)randomly, and shares the green list ${\mathcal{G}}$ with the detector as part of the watermark key. The watermark is embedded by prompting the use of green tokens during decoding: given the NTP distribution ${\bm{P}}_{t}$ and the green list ${\mathcal{G}}$, the watermarked decoder constructs an alternative distribution ${\bm{Q}}_{t}=(Q_{t,w})_{w\in{\mathcal{W}}}$ and samples the next token from ${\bm{Q}}_{t}$. The simplest version is the “hard” watermark, where each sampled token is limited in the green list ${\mathcal{G}}$:
$Q_{t,w}=\begin{cases}P_{t,w}/P_{t,{\mathcal{G}}}\quad&\text{if }w\in{\mathcal{G}}\\ 0&\text{otherwise}\end{cases}$ (1)
@@
To aid detection, we share the information $\xi_{t}=({\mathcal{G}},\zeta_{t})$, which includes the green list ${\mathcal{G}}$ and the random variable $\zeta_{t}$, with the detector. This scheme benefits from the following property of $\zeta_{t}$ as a result of the maximal coupling process, and the proof is provided in Section 3.1:
@@
We utilize the conditional distribution of $\zeta_{t}$ given the information $({\bm{P}}_{t},{\mathcal{G}})$ to detect the existence of watermarks via a hypothesis testing framework introduced in (Li et al., 2024), which lays two fundamental working hypotheses. In the null hypothesis $H_{0}$, this framework assumes that a human writes the text without the shared watermark key $({\mathcal{G}},\zeta_{t})$. In this way, the computed $\zeta_{t}$ doesn’t correlate with the generation of the given text so the identified green tokens should be independent of each $\zeta_{t}$, i.e., each $\zeta_{t}$ acts as an i.i.d. copy from $U[0,1]$, regardless of $\bm{P}_{t}$ and ${\mathcal{G}}$. Under the alternative hypothesis $H_{1}$, the framework assumes that the text is generated using the specified watermarked decoder with the pseudorandom variables $\zeta_{1:n}$. Consequently, $\zeta_{t}$ has a known conditional distribution, as described in Lemma 1.1. This reduces the watermark detection problem to identifying distributional differences in the values of $\zeta_{t}$.
@@ 1. $H_{1}^{(\mathrm{mix})}$:
    Without sufficient watermark signals, i.e., $p+2q>1$, $H_{0}$ and $H_{1}^{(\mathrm{mix})}$ merge asymptotically. As a result, no test can reliably separate $H_{0}$ and $H_{1}^{(\mathrm{mix})}$.
@@
Another application of maximal coupling is speculative decoding (Leviathan et al., 2023; Chen et al., 2023), where a smaller draft model generates speculative token predictions that are then accepted or corrected by a more powerful target model. We interpret this setup as a model of human-language model interaction, where the smaller model represents the watermarked language model’s text generation capabilities, and the target model represents the human’s attempt to refine the generated text, focusing solely on enhancing the overall text quality. In Section 5, we apply this setup to our watermarking scheme and demonstrate that our method effectively retains the watermark signal despite the targeted attempts to modify the text. Moreover, compared to Aaronson and Kirchner (2023)’s approach, our method results in fewer text modifications, suggesting that users would likely need to invest less editing effort to achieve their desired text quality when working with our watermarking scheme.
Our Contributions.
We introduce a novel unbiased variant of the green/red list watermark by leveraging maximal coupling. After formulating the robust watermark detection problem as a sparse detection problem, we propose several detection methods and explore its theoretical properties. In particular, we provide theoretical insights into the statistical power of these detection methods, corroborated by extensive simulation studies. We demonstrate the effectiveness of our watermarking scheme against multiple baselines via experiments on large language models of varying sizes, as well as under a targeted text modification setup. Our findings offer a promising direction for watermarking in language models, balancing the need for detectability with minimal impact on text quality.
@@ 2 Preliminaries
Unbiased Watermarking Schemes.
Kirchenbauer et al. (2023a) introduced a “soft” watermark to mitigate the bias of the “hard” watermark by sampling from both the green list ${\mathcal{G}}$ and the red list ${\mathcal{R}}:={\mathcal{W}}\setminus{\mathcal{G}}$, but skewing the sampling distribution to favor tokens from the green list. That is, given a constant $\delta>0$ and the next token distribution $P_{w}$, we construct the alternative distribution $Q_{w}$ as $Q_{w}=(\mathrm{e}^{\delta}P_{w})/C$ if $w\in{\mathcal{G}}$ and $Q_{w}=P_{w}/C$ if $w\not\in{\mathcal{G}}$, where $C=\mathrm{e}^{\delta}P_{{\mathcal{G}}}+P_{{\mathcal{W}}\setminus{\mathcal{G}}}=1+(\mathrm{e}^{\delta}-1)P_{{\mathcal{G}}}$. Both watermarks are biased according to the following definition (Kuditipudi et al., 2023; Li et al., 2024). [^5] In this setup, ${\mathcal{S}}$ as a function of the next token distribution $\bm{P}$ and a random variable $\xi={\mathcal{G}}$, the (pseudo)randomly generated green list. For the proof of the biasedness, we refer the reader to Appendix A.1.
@@
The method by Aaronson and Kirchner (2023) adapts exponential minimal sampling and achieves unbiasedness when the random variables $\xi_{t}$ are truly i.i.d. However, in practice, the pseudorandom variable $\xi_{t}$ is deterministically dependent on the prior $k$ tokens $w_{(t-k)},...,w_{(t-1)}$. This can lead to repetitive texts and quality degradation, particularly when the same $k$ tokens appears multiple times in the generated text (Christ et al., 2023; Kuditipudi et al., 2023). While Hu et al. (2023); Wu et al. (2023); Dathathri et al. (2024) propose repeated context masking, a method that prevents applying watermark on step $t$ if the prior $k$ tokens have been used to watermark previously, our experiments in Section 5 demonstrate that this decoder still produces text that is more repetitive and lower quality compared to other watermarked decoders. In contrast, our proposed stochastic decoder incorporates additional randomness for sampling and the pseudorandom variable $\zeta_{t}$ is only used for rejection. Several theoretical analyses and variants of the deterministic-decoder-based method have been proposed (Fernandez et al., 2023; Zhao et al., 2024), but no existing work completely addresses this issue. Another method is the inverse transform sampling strategy (Christ et al., 2023; Kuditipudi et al., 2023; Hu et al., 2023; Li et al., 2024). However, these methods typically yield weaker power than the exponential minimal sampling method, even when true randomness is applied.
Building on inverse transform sampling, Wu et al. (2023) proposed an unbiased watermarking scheme that incorporates random permutation and a probability reweighting strategy: given a pseudorandom permutation of the vocabulary ${\mathcal{W}}$ and a parameter $\alpha<0.5$, this scheme samples from the distribution $\bm{Q}$, where $Q_{w^{(i)}}=F_{w^{(i)}}-F_{w^{(i-1)}}$ and $F_{w^{(i)}}=\max\left\{\sum_{j=1}^{i}P_{w^{(j)}}-\alpha,0\right\}+\max\left\{\sum_{j=1}^{i}P_{w^{(j)}}-(1-\alpha),0\right\}.$ Here, $w^{(i)}$ denotes the $i$th token in the permutation, with the convention that $F(w^{(0)})=0$. While this method is unbiased, tuning the hyperparameter $\alpha$ can be a delicate task – a limitation that our proposed method eliminates. Furthermore, as shown in Section 5, our method achieves higher detection power. For a more extensive review of unbiased watermarking schemes, see Appendix A.1.
Maximal Coupling.
Maximal Coupling is a powerful tool in probability theory that constructs a coupling between two random variables to maximize the probability that they are equal (Thorisson, 2000; Roch, 2024). In our watermarking scheme, we use maximal coupling to sample tokens in a way that preserves the original next-token distribution $\bm{P}_{t}$ while incorporating information from an alternative distribution $\bm{Q}_{t}$, both supported on the vocabulary ${\mathcal{W}}$. Algorithm 1 outlines the procedure to sample $w$ from the NTP distribution $\bm{P}$, where both $\min$ and $\max$ operations are applied element-wise.
@@ 3 Watermarking Scheme via Maximal Coupling
We formulate the watermarking procedure within a communication protocol involving three parties: the language model provider, the user, and the detector. The protocol for embedding and detecting a watermark in text generated by the language model is outlined in Algorithm 2.
@@
This protocol implies that a watermarking scheme has two parts: (1) a decoding (or sampling) algorithm that can embed identifiable statistical signals into the generated text, and (2) a detection algorithm that can detect the watermark signal in the received text. In this remainder of this section, we introduce every aspect of our proposed watermark, including different detection methods, asymptotic theoretical analysis on detection boundaries, and practical considerations.
Setup.
For clarity in illustrating the watermarking scheme, we first describe a simplified setup for the communication protocol and delay the adaptation in practice to Section 3.3. In particular, we assume that: (1) the text generation has a predefined length $n$; (2) the language model provider randomly generates the watermark key $\bm{\xi}=({\mathcal{G}}_{1:n},\zeta_{1:n})$ and shares it with the detector: ${\mathcal{G}}_{1:n}=({\mathcal{G}}_{1},\ldots,{\mathcal{G}}_{n})$ is a set of green lists, and $\zeta_{1:n}=(\zeta_{1},...,\zeta_{n})$ consists of independent standard uniform random variables; (3) the language model provider uses ${\mathcal{G}}_{1:n}$ and $\zeta_{1:n}$ to generate a single output text $w_{1:n}=(w_{1},...,w_{n})$ given an arbitrary prompt ${\mathcal{P}}$; and (4) the user only modifies the text via substitution attack, i.e., the user replaces some tokens in the generated text with other tokens so that $l=n$ in this protocol.
@@ 3.1 Decoding
     Define $\bm{Q}_{t}$ as in Equation (1), i.e. $Q_{t,w}=P_{t,w}\mathds{1}_{w\in{\mathcal{G}}_{t}}/P_{t,{\mathcal{G}}_{t}}$.
@@
The decoding algorithm works as follows: for $t=1,...,n$, given the token distribution $\bm{P}_{t}$ at step $t$, we first construct an alternative distribution $\bm{Q}_{t}$ as in Equation (1). Then we use $\zeta_{t}$ to sample the next token $w$ via maximal coupling. This process is summarized in Algorithm 3. As an edge case, if $P_{t,{\mathcal{G}}_{t}}=0$, we simply sample from $\bm{P}_{t}$. Note that this decoding scheme is unbiased by Lemma 2.2. As $\min(P_{t,w},Q_{t,w})=P_{t,w}\mathds{1}_{w\in{\mathcal{G}}_{t}}$ and $\max(0,P_{t,w}-Q_{t,w})=P_{t,w}\mathds{1}_{w\not\in{\mathcal{G}}_{t}}$, our watermarked decoder is effectively a stochastic function ${\mathcal{S}}(\bm{P}_{t},{\mathcal{G}}_{t},\zeta_{t})$ sampling from the following distribution:
$\mathbb{P}[w_{t}=w|\bm{P}_{t},{\mathcal{G}}_{t},\zeta_{t}]=\mathds{1}_{\zeta_{t}\leq P_{t,{\mathcal{G}}_{t}}}\frac{P_{t,w}\mathds{1}_{w\in{\mathcal{G}}_{t}}}{P_{t,{\mathcal{G}}_{t}}}+\mathds{1}_{\zeta_{t}>P_{t,{\mathcal{G}}_{t}}}\frac{P_{t,w}\mathds{1}_{w\not\in{\mathcal{G}}_{t}}}{1-P_{t,{\mathcal{G}}_{t}}},$

### v2

@@ 1 Introduction
The field of artificial intelligence (AI) has seen rapid growth in recent years, with large language models (LLMs) playing a central role in many applications. These models, such as OpenAI’s ChatGPT (Achiam et al., 2023), are capable of understanding and generating human-like text. However, the potential for misuse has raised concerns about the trustworthiness of AI-generated content, particularly in the context of disinformation, fraud, social engineering, and cybercrime (Europol, 2023). To mitigate these risks, President Biden issued an executive order on Safe, Secure, and Trustworthy Artificial Intelligence, which emphasized the need for standards to detect AI-generated content and authenticate official content, setting a precedent for transparency and trust in digital communications (Biden, 2023). In particular, the Department of Commerce identified watermarking as a pivotal means to clearly label AI-generated content.
@@
The earliest watermarking scheme, the Gumbel-max watermark, was introduced by Aaronson and Kirchner (2023). This method selects the next token $w_{t}$ using the decoder $w_{t}={\mathcal{S}}(\bm{P}_{t},\xi_{t}):=\operatorname{argmax}_{w\in{\mathcal{W}}}\log\frac{U_{w}}{P_{w}}$, assuming $\xi_{t}=(U_{1},\ldots,U_{|{\mathcal{W}}|})$ is a random vector consisting of $|{\mathcal{W}}|$ i.i.d. copies of $U[0,1]$, serving as part of the watermark key. While this watermarked decoder theoretically preserves the original next-token distribution when $\xi_{t}$ is truly random and achieves high detection efficiency, this decoder can generate repetitive text and significantly degrade text quality in practice (Kuditipudi et al., 2023). The main reason is that, for existing implementations, the pseudorandom variable $\xi_{t}$ depends deterministically on the prior $k$ tokens $w_{(t-k)},\ldots,w_{(t-1)}$. As a result, when conditioned on previous tokens, the decoder ${\mathcal{S}}$ deterministically outputs the next token, which can lead to sub-optimal text generation (Holtzman et al., 2020).
Another influential watermarking scheme, the green/red list watermark, employs a stochastic decoder. First introduced by Kirchenbauer et al. (2023a), this approach has gained popularity (Fernandez et al., 2023; Wu et al., 2023). It first partitions the token vocabulary ${\mathcal{W}}$ into the green list ${\mathcal{G}}$ and the red list ${\mathcal{R}}={\mathcal{W}}\setminus{\mathcal{G}}$ (pseudo)randomly, and shares the green list ${\mathcal{G}}$ with the detector as part of the watermark key. The watermark is embedded by prompting the use of green tokens during decoding: given the NTP distribution ${\bm{P}}_{t}$ and the green list ${\mathcal{G}}$, the watermarked decoder constructs an alternative distribution ${\bm{Q}}_{t}=(Q_{t,w})_{w\in{\mathcal{W}}}$ and samples the next token from ${\bm{Q}}_{t}$. The simplest version is the “hard” watermark, where each sampled token is limited in the green list ${\mathcal{G}}$:
$Q_{t,w}=\frac{P_{t,w}}{P_{t,{\mathcal{G}}}}\mathbb{1}_{\{w\in{\mathcal{G}}\}},$ (1)
@@
To aid detection, we share the information $\xi_{t}=({\mathcal{G}},\zeta_{t})$, the green list ${\mathcal{G}}$ and the random variable $\zeta_{t}$, with the detector. This scheme benefits from the following property of $\zeta_{t}$ as a result of the maximal coupling process, and the proof is provided in Section 3.1:
@@
We utilize the conditional distribution of $\zeta_{t}$ given $({\bm{P}}_{t},{\mathcal{G}})$ to detect the existence of watermarks via a hypothesis testing framework introduced by Li et al. (2024), which lays two fundamental working hypotheses. Under the null hypothesis $H_{0}$, this framework assumes that a human writes the text without the shared watermark key $({\mathcal{G}},\zeta_{t})$. The computed $\zeta_{t}$ doesn’t correlate with the generation of the given text so the identified green tokens should be independent of each $\zeta_{t}$, i.e., each $\zeta_{t}$ acts as an i.i.d. copy from $U[0,1]$, regardless of $\bm{P}_{t}$ and ${\mathcal{G}}$. Under the alternative hypothesis $H_{1}$, the text is assumed to be generated using the specified watermarked decoder with the pseudorandom variables $\zeta_{1:n}$ Consequently, $\zeta_{t}$ has a known conditional distribution, as described in Lemma 1.1. This reduces the watermark detection problem to identifying distributional differences in the values of $\zeta_{t}$.
@@ 1. $H_{1}^{(\mathrm{mix})}$:
    Without sufficient watermark signals, i.e., $p+2q>1$, $H_{0}$ and $H_{1}^{(\mathrm{mix})}$ merge asymptotically. Hence, no test can reliably separate $H_{0}$ and $H_{1}^{(\mathrm{mix})}$.
@@
Another application of maximal coupling is speculative decoding (Leviathan et al., 2023; Chen et al., 2023), where a smaller draft model generates speculative token predictions that are then accepted or corrected by a more powerful target model. We interpret this setup as a model of human-LLM interaction, where the smaller model represents the watermarked language model’s text generation capabilities, and the target model represents the human’s attempt to refine the generated text, focusing solely on enhancing the overall text quality. In Section 5, we apply this setup to our watermarking scheme and demonstrate that our method effectively retains the watermark signal despite the targeted attempts to modify the text. Compared to Aaronson and Kirchner (2023)’s approach, our method results in fewer text modifications, suggesting that users would likely need to invest less editing effort to achieve their desired text quality when working with our watermarking scheme.
Our Contributions. We introduce a novel unbiased variant of the green/red list watermark by leveraging maximal coupling. After formulating the robust watermark detection problem as a sparse detection problem, we propose several detection methods and explore its theoretical properties. In particular, we provide theoretical insights into the statistical power of these detection methods, corroborated by extensive simulation studies. We demonstrate the effectiveness of our watermarking scheme against multiple baselines via experiments on large language models of varying sizes, as well as under a targeted text modification setup. Our findings offer a promising direction for watermarking in language models, balancing the need for detectability with minimal impact on text quality.
@@ 2 Preliminaries
Unbiased Watermarking Schemes. Kirchenbauer et al. (2023a) introduced a “soft” watermark to mitigate the bias of the “hard” watermark by sampling from both the green list ${\mathcal{G}}$ and the red list ${\mathcal{R}}:={\mathcal{W}}\setminus{\mathcal{G}}$, but skewing the sampling distribution to favor tokens from the green list. That is, given a constant $\delta>0$ and the next token distribution $P_{w}$, we construct the alternative distribution $Q_{w}$ as $Q_{w}=(\mathrm{e}^{\delta}P_{w})/C$ if $w\in{\mathcal{G}}$ and $Q_{w}=P_{w}/C$ if $w\not\in{\mathcal{G}}$, where $C=\mathrm{e}^{\delta}P_{{\mathcal{G}}}+P_{{\mathcal{W}}\setminus{\mathcal{G}}}=1+(\mathrm{e}^{\delta}-1)P_{{\mathcal{G}}}$. Both watermarks are biased according to the following definition (Kuditipudi et al., 2023; Li et al., 2024). [^5] In this setup, ${\mathcal{S}}$ as a function of the next token distribution $\bm{P}$ and a random variable $\xi={\mathcal{G}}$, the (pseudo)randomly generated green list. For the proof of the biasedness, we refer the reader to the online Appendix.
@@
Aaronson and Kirchner (2023) adapts exponential minimal sampling and achieves unbiasedness when the random variables $\xi_{t}$ are truly i.i.d. However, in practice, the pseudorandom variable $\xi_{t}$ is deterministically dependent on the prior $k$ tokens $w_{(t-k)},...,w_{(t-1)}$. This can lead to repetitive texts and quality degradation, particularly when the same $k$ tokens appears multiple times in the generated text (Kuditipudi et al., 2023). While Hu et al. (2023); Dathathri et al. (2024) propose repeated context masking to prevent applying watermark on step $t$ if the prior $k$ tokens have been used to watermark previously, our experiments in Section 5 demonstrate that this decoder still produces text that is more repetitive and lower quality compared to other watermarked decoders. Several theoretical analyses and variants of the deterministic-decoder-based method have been proposed (Fernandez et al., 2023; Zhao et al., 2024), but no existing work completely addresses this issue. In contrast, our proposed stochastic decoder incorporates additional randomness for sampling and the pseudorandom variable $\zeta_{t}$ is only used for rejection. Another method is the inverse transform sampling strategy (Christ et al., 2024; Kuditipudi et al., 2023; Hu et al., 2023; Li et al., 2024). However, these methods typically yield weaker power than the exponential minimal sampling. Building on inverse transform sampling, Wu et al. (2023) proposed an unbiased watermarking scheme that incorporates random permutation and a probability reweighting strategy. [^7] Given a pseudorandom permutation of the vocabulary ${\mathcal{W}}$ and a parameter $\alpha<0.5$, this scheme samples from the distribution $\bm{Q}$, where $Q_{w^{(i)}}=F_{w^{(i)}}-F_{w^{(i-1)}}$ and $F_{w^{(i)}}=\max\left\{\sum_{j=1}^{i}P_{w^{(j)}}-\alpha,0\right\}+\max\left\{\sum_{j=1}^{i}P_{w^{(j)}}-(1-\alpha),0\right\}.$ Here, $w^{(i)}$ denotes the $i$th token in the permutation, with the convention that $F(w^{(0)})=0$. While this method is unbiased, tuning the hyperparameter can be a delicate task – a limitation that our proposed method eliminates. Furthermore, as shown in Section 5, our method achieves higher detection power. For a more extensive review of unbiased watermarking schemes, see the online Appendix.
@@
Maximal Coupling. Maximal coupling is a powerful tool in probability theory that constructs a coupling between two random variables to maximize the probability that they are equal (Thorisson, 2000). We use maximal coupling to sample tokens in a way that preserves the original next-token distribution $\bm{P}_{t}$ while incorporating information from an alternative distribution $\bm{Q}_{t}$, both supported on the vocabulary ${\mathcal{W}}$. Algorithm 1 outlines the procedure to sample $w$ from the NTP distribution $\bm{P}$, where both $\min$ and $\max$ operations are applied element-wise, and the proof for Lemma 2.2 can be found in the online Appendix.
@@ 3 Watermarking Scheme via Maximal Coupling
We formulate the watermarking procedure within a communication protocol involving three parties: the language model provider, the user, and the detector. As outlined in Algorithm 2, this protocol implies that a watermarking scheme has two parts: (1) a decoding (or sampling) algorithm that can embed identifiable statistical signals into the generated text, and (2) a detection algorithm that can detect the watermark signal in the received text. In this remainder of this section, we introduce every aspect of our proposed watermark, including different detection methods, asymptotic theoretical analysis on detection boundaries, and practical considerations.
@@
Setup. For clarity in illustrating the watermarking scheme, we first describe a simplified setup for the communication protocol and delay the adaptation in practice to Section 3.3. In particular, we assume that: (1) the text generation has a predefined length $n$; (2) the language model provider randomly generates the watermark key $\bm{\xi}=({\mathcal{G}}_{1:n},\zeta_{1:n})$ and shares it with the detector: ${\mathcal{G}}_{1:n}=({\mathcal{G}}_{1},\ldots,{\mathcal{G}}_{n})$ is a set of green lists, and $\zeta_{1:n}=(\zeta_{1},...,\zeta_{n})$ consists of independent standard uniform random variables; (3) the language model provider uses ${\mathcal{G}}_{1:n}$ and $\zeta_{1:n}$ to generate a single output text $w_{1:n}=(w_{1},...,w_{n})$ given an arbitrary prompt ${\mathcal{P}}$; and (4) the user only modifies the text via substitution attack, i.e., the user replaces some tokens in the generated text with other tokens so that $l=n$ in this protocol.
@@ 3.1 Decoding
     Define $\bm{Q}_{t}$ as in Equation (1), i.e. $Q_{t,w}=P_{t,w}\mathbb{1}_{w\in{\mathcal{G}}_{t}}/P_{t,{\mathcal{G}}_{t}}$.
@@
The decoding algorithm works as follows: for $t=1,...,n$, given the token distribution $\bm{P}_{t}$ at step $t$, we first construct an alternative distribution $\bm{Q}_{t}$ as in Equation (1). Then we use $\zeta_{t}$ to sample the next token $w$ via maximal coupling. This process is summarized in Algorithm 3. As an edge case, if $P_{t,{\mathcal{G}}_{t}}=0$, we simply sample from $\bm{P}_{t}$. Note that this decoding scheme is unbiased by Lemma 2.2. As $\min(P_{t,w},Q_{t,w})=P_{t,w}\mathbb{1}_{w\in{\mathcal{G}}_{t}}$ and $\max(0,P_{t,w}-Q_{t,w})=P_{t,w}\mathbb{1}_{w\not\in{\mathcal{G}}_{t}}$, our watermarked decoder is effectively a stochastic function ${\mathcal{S}}(\bm{P}_{t},{\mathcal{G}}_{t},\zeta_{t})$ sampling from the following distribution:
$\mathbb{P}[w_{t}=w|\bm{P}_{t},{\mathcal{G}}_{t},\zeta_{t}]=\mathbb{1}_{\zeta_{t}\leq P_{t,{\mathcal{G}}_{t}}}\frac{P_{t,w}\mathbb{1}_{w\in{\mathcal{G}}_{t}}}{P_{t,{\mathcal{G}}_{t}}}+\mathbb{1}_{\zeta_{t}>P_{t,{\mathcal{G}}_{t}}}\frac{P_{t,w}\mathbb{1}_{w\not\in{\mathcal{G}}_{t}}}{1-P_{t,{\mathcal{G}}_{t}}},$
@@
Remark 3.3.
Our decoder can also be viewed as an extension of that proposed by Christ et al. (2024), which developed a watermarking technique for binary language models ($\mathcal{M}^{(b)}$) where the vocabulary is limited to two tokens ($\mathcal{W}^{(b)}=\{0,1\}$). In this binary context, the watermark key $\zeta$ in Algorithm 1 directly determines whether the next sampled token is 00 or $1$, making our method functionally equivalent to their proposal. To adapt this binary model to accommodate larger vocabularies ($|\mathcal{W}|\gg 2$), Christ et al. (2024) suggested representing each token as a $\log|\mathcal{W}|$-bit string and applying binary language models to decode these strings. However, this adaptation lacks practicality since language models operate natively with large vocabularies. In contrast, our approach avoids the extra step of manually decomposing token probabilities into bits by directly leveraging the model’s inherent probability distribution. Our method provides a more natural extension by working directly with these large vocabulary sets rather than requiring binary encoding and decoding steps, by conceptualizing all green tokens as one category and all red tokens as another.